# Powerformer: A Section-adaptive Transformer for Power Flow Adjustment

### v2

@@
Abstract
In this paper, we present a novel transformer architecture tailored for learning robust power system state representations, which strives to optimize power dispatch for the power flow adjustment across different transmission sections. Specifically, our proposed approach, named Powerformer, develops a dedicated section-adaptive attention mechanism, separating itself from the self-attention used in conventional transformers. This mechanism effectively integrates power system states with transmission section information, which facilitates the development of robust state representations. Furthermore, by considering the graph topology of power system and the electrical attributes of bus nodes, we introduce two customized strategies to further enhance the expressiveness: graph neural network propagation and multi-factor attention mechanism. Extensive evaluations are conducted on three power system scenarios, including the IEEE 118-bus system, a realistic 300-bus system in China, and a large-scale European system with 9241 buses, where Powerformer demonstrates its superior performance over several baseline methods.
keywords:
@@ Abstract
journal: xxx
Chen et al.: Powerformer: A Section-adaptive Transformer for Power Flow Adjustment
Graph Abstract:
@@
1 Introduction
In real-world scenarios, power system operators typically focus on monitoring the operational state of a power system that involves numerous variables, placing emphasis on transmission sections rather than isolated components. As shown in Fig. 1, a transmission section consists of a group of transmission lines with the same color. To ensure power system security margins, operators commonly employ the total transfer capability of transmission sections to monitor power flows and regulate the operational state de2024discrete ; avraam2023operational ; jiang2022steady . Specifically, overloading a critical transmission section poses a serious threat to the system, potentially triggering a chain reaction of failures that may cause severe cascading blackouts. Hence, the efficient management of power flow at transmission sections is paramount, serving as a fundamental pillar to ensure the stability, resilience, and reliability of the power system.
Figure 1: The illustration of 10 transmission sections $\{\Phi_{1},...,\Phi_{10}\}$ in the IEEE 118-case system, where each section is represented by a set of transmitting lines that share the same color and are closely located electrically. An illustrative section $\Phi_{7}$ is the set of purple lines, i.e., $\Phi_{7}=\,${bus33-bus37, bus19-bus34, bus30-bus38, bus23-bus24}, that partitions the system into distinct left and right components.
To address power flow adjustment across different transmission sections, power dispatch is an effective strategy that enables precise finetuning of generation outputs to optimize power flow distribution; traditional dispatch methods, which rely on the system model, can be primarily categorized into two distinct classes: optimization-based capitanescu2011state ; qiang2023enhancing and sensitivity-based dutta2008optimal ; lai2021vulnerability techniques. Optimization-based methods utilize numerical techniques to transform the dispatch problem into a constrained programming problem, while sensitivity-based methods apply iterative calculations to find candidate generators by calculating sensitivity indices. However, these model-based methods often encounter the computational burdens and struggle to effectively control unpredictable system coupling within a limited time. Thus, traditional methods are unsuitable for effectively managing real-time control in a complex and large-scale power system.
Recently, deep reinforcement learning (DRL) has emerged as an innovative alternative to traditional power dispatch methods. This data-driven approach effectively utilizes the robust capabilities of the pertained deep neural networks, providing a powerful solution for various power flow adjustment. However, among the existing DRL methods li2023deep ; tang2023energy ; deng2024novel ; liu2023ask , there are two evident limitations in the state representations: 1) the absence of efficient awareness regarding transmission sections, which hinders the optimal performance for distinct section adjustment; 2) the high coupled electrical factors and the ignored topology structure of power system, ignificantly affect the extraction of the state characteristics.
In this paper, we strive to develop an advanced transformer architecture for learning robust power system state representations, facilitating precise power dispatch strategies that stabilize the power system under various section settings. To tackle the first limitation, we develop a section-adaptive attention mechanism to enable an efficient amalgamation of section information with the power system state, leveraging the comprehensive capabilities of the transformer architecture vaswani2017attention ; zhang2023knowledge ; zhang2022bootstrapping for the specific sectional system analysis. For the second limitation, we begin by analyzing the graph topology of the power system, and employing graph neural networks xu2018powerful ; Li2023MessagepassingST ; VERDONE2024122151 for feature propagation to embed structural information into state representations. Furthermore, we introduce a multi-factor attention mechanism to efficiently fuse each state factor, ensuring comprehensive and accurate extraction of system states. Finally, we utilize the proposed Powerformer, integrating the training of Dueling DQN Wang2015DuelingNA to achieve precise and efficient control of power flow across transmission sections, thereby leading to improvements in the overall system performance.
Our main contributions can be summarized as follows:
  • 1.
  • 2.
  • 3.
    The extensive evaluation on three distinct power system cases, including the IEEE 118-bus system, a realistic 300-bus system in China, and a large-scale 9241-bus European system, demonstrates our method’s superiority over several baselines with significantly improved performance.
2 PRELIMINARY
2.1 Problem Formulation and Task Setting
2.1.1 Problem Formulation
Transmission section power flow adjustment involves the regulation and control of electrical power flow through the specific section or interface of a power transmission network. In a specific power system, there exist $N$ transmission sections $\{\Phi_{1},\Phi_{2},...,\Phi_{N}\}$, where each section $\Phi_{i}$ consists of several transmission lines that not only share the same direction of active power flow but also maintain close electrical proximity to one another liu2023transmission . To accurately determine the total power flow of a transmission section $\Phi_{i}$, the combined power contributions are aggregated and meticulously calculated:
$\mathcal{P}_{\Phi_{i}}=\sum_{j\in\Phi_{i}}\mathcal{P}_{j}\quad\,{s.t.}\;\mathcal{P}_{\text{min},\Phi_{i}}\leq\,\mathcal{P}_{\Phi_{i}}\,\leq\mathcal{P}_{\text{max},\Phi_{i}},$ (1)
where $\mathcal{P}_{\Phi_{i}}$ and $\mathcal{Q}_{\Phi_{i}}$ represent the total active and reactive power flow in the $i$-th transmission section, and $\mathcal{P}_{j}$ and $\mathcal{Q}_{j}$ denote the active and reactive power flow in the $j$-th transmission line within the section $\Phi_{i}$. $\mathcal{P}_{\text{min},\Phi_{i}}$, $\mathcal{P}_{\text{max},\Phi_{i}}$, $\mathcal{Q}_{\text{min},\Phi_{i}}$, $\mathcal{Q}_{\text{max},\Phi_{i}}$ are the minimum and maximum allowable active and reactive power flow in the $i$-th transmission section. These limits are essential for ensuring the power system operates safely and efficiently, maintaining power flow within each transmission section at levels that prevent overloading and stability issues. To address this problem, power dispatch is a commonly employed strategy for regulating power system flow and maintaining safety margin constraints.
2.1.2 Task Setting
In the context of power flow adjustment across various transmission sections, multi-task learning zhang2018overview ; TAN2023121177 is suitable to manage power dispatch for simultaneously addresses different scenarios within the power system, where each section’s power flow treated as a distinct task managed by a unique adjustment strategy. In this study, we establish the 10-section power flow adjustment task, selected based on the modularity of communities to reflect the importance of the sections, in order to evaluate the model’s performance on three distinct power system cases.
2.2 Sectional Power System Network
The power system topology, being non-Euclidean chen2020covariance ; Chen2022RiemannianLM , is typically represented as a graph structure chen2023hybrid ; sun2022deep ; chen2023tele ; chen2023improving . Therefore, we organize power system data in the form of graph-structured data. In a power system graph network, nodes represent buses and edges symbolize transmission lines. Specifically, a power system network with section information can be defined as:
Definition 1 (Sectional Power System Network) A sectional power system network is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E},\boldsymbol{H},\boldsymbol{Z}_{\Phi})$, where $\mathcal{V}=\{v_{1},...,v_{n}\}$ and =$\mathcal{E}=\{e_{1},...,e_{m}\}$ represent the sets of $n$ nodes and $m$ edges, respectively. $\boldsymbol{H}=[\boldsymbol{h}_{1},...,\boldsymbol{h}_{n}]\in\mathbb{R}^{d\times n}$ denotes the state feature matrix, where $\boldsymbol{h}_{i}$ denotes the $i$-th node. $\boldsymbol{Z}_{\Phi}$ is the representation of transmission section $\Phi$.
To illustrate a sectioned power system network, Fig. 1 provides an example with 10 sections in the IEEE 118-case system. In this case, transmission section $\Phi_{7}$ is represented by the purple lines, namely $\Phi_{7}=${bus33-bus37, bus19-bus34, bus30-bus38, bus23-bus24}. This specific section divides the system into left and right components.
2.3 Transformer Architecture
Numerous Pre-trained Language Models (PLMs) employ a multi-layer Transformer architecture vaswani2017attention for the purpose of encoding text representations. In each layer of the Transformer, a multi-head self-attention mechanism is utilized to generate contextualized representations for individual text tokens. Let $\boldsymbol{H}^{(l)}=[\boldsymbol{h}_{1}^{(l)},...,\boldsymbol{h}_{n}^{(l)}]$ denote the output feature of the $l$-th Transformer layer, where $\boldsymbol{h}_{i}^{(l)}\in\mathbb{R}^{d}$ is the hidden representation of the text token at position $i$. Subsequently, in the $(l+1)$-th Transformer layer, the multi-head self-attention (MHA) is calculated as follows:
@@ 2.3 Transformer Architecture
$\text{head}^{t}(\boldsymbol{H}^{(l)})=\boldsymbol{V}_{t}^{(l)}\cdot\text{softmax}\left(\frac{{\boldsymbol{K}_{t}^{(l)\top}\boldsymbol{Q}_{t}^{(l)}}}{\sqrt{d}}\right),$ (3)
$\boldsymbol{Q}_{t}^{(l)}=\boldsymbol{W}_{Q,t}^{(l)}\boldsymbol{H}^{(l)},\;\boldsymbol{K}_{t}^{(l)}=\boldsymbol{W}_{K,t}^{(l)}\boldsymbol{H}^{(l)},\;\boldsymbol{V}_{t}^{(l)}=\boldsymbol{W}_{V,t}^{(l)}\boldsymbol{H}^{(l)},$ (4)
[…]
3 Method
3.1 Overall Architecture
An overview of the Powerformer architecture is presented in Fig. 2b, illustrating the fundamental components crucial for obtaining robust representations in sectional power system networks, such as state factorization, Graph Neural Network (GNN) propagation, and section-adaptive attention, etc. Different from the conventional Transformer, our architecture utilizes the set $\{\boldsymbol{H},\boldsymbol{Z}_{\Phi},\mathcal{E}\}$ as input, which consists of the state feature matrix, section power flow, and graph structure information. The section-specific information $\boldsymbol{Z}_{\Phi}$ is utilized to learn Query representation, enabling the development of section-adaptive attention mechanism. To address the specific electrical factors on buses, we disentangle the system state representation into $k$ distinct factors to prevent feature coupling during propagation, and then incorporate the section-adaptive attention matrices to establish the multi-factor section-adaptive attention mechanism. Furthermore, to enhance the expressiveness of Key, Value propagation, we utilize the popular GNN architecture to embed the graph structure information $\mathcal{E}$ into the final state representation.
Figure 2: The illustration of the Transformer and Powerformer architectures. (a) Transformer architecture with the multi-head self-attention mechanism (MHA). (b) Powerformer architecture with the multi-factor section-adaptive attention mechanism (MFSA).
3.2 Powerformer for Sectional Power System
If the state matrix is directly inputted into the propagation layer, it will result in the coupling of different factors, causing the redundancy in the learned features and ultimately diminishing the expressive power. Thus, we begin by disentangling the input state feature into $k=4$ factors, with the $l$-th layer’s representation of the $t$-th factor denoted by $\boldsymbol{H}_{t}^{(l)}\in\mathbb{R}^{d\times n}$.
$\boldsymbol{K}_{t}^{(l)}=\mathcal{F}_{K,t}^{(l)}(\boldsymbol{H}_{t}^{(l)},\mathcal{E}),\;\boldsymbol{V}_{t}^{(l)}=\mathcal{F}_{V,t}^{(l)}(\boldsymbol{H}_{t}^{(l)},\mathcal{E}),$ (5)
where $\mathcal{F}_{K,t}^{(l)},\mathcal{F}_{V,t}^{(l)}$ represent the GNN propagation to update the Key and Value: $\boldsymbol{K}_{t}^{(l)},\boldsymbol{V}_{t}^{(l)}\in\mathbb{R}^{d\times n}$. Specifically, this propagation is implemented using a Graph Isomorphism Network (GIN) model xu2018powerful :
$\boldsymbol{h}_{v}^{(l)}={\rm MLP}^{(l)}((1+\epsilon^{(l)})\cdot\boldsymbol{h}_{v}^{(l-1)}+\sum\nolimits_{u\in\mathcal{N}(v)}\boldsymbol{h}_{u}^{(l-1)}),$ (6)
where $\boldsymbol{h}_{v}^{(l)}$ represents the feature vector of node $v$ at the $l$-th layer. $\mathcal{N}(v)$ refers to the neighborhood set of node $v$ based on the edge set $\mathcal{E}$. MLP is the Multi-layer Perceptron network.
Section-adaptive Attention (SA). For section-adaptive representations, we first utilize the power flow information of the transmission lines on the current section to encode the section, and set the power flow information that is not within the section to zero. The resulting initialization representation is $\boldsymbol{Z}_{\Phi}$. Then, we use a MLP network to learn the Query component and align its dimensions with $\boldsymbol{K}_{t}^{(l)}$ and $\boldsymbol{V}_{t}^{(l)}$:
$\boldsymbol{Q}^{(l)}=\mathcal{M}_{Q}^{(l)}(\boldsymbol{Z}_{\Phi}^{(l)}),$ (7)
where $\mathcal{M}_{Q}^{(l)}$ is the employed MLP to update section representation. The updated Query is $\boldsymbol{Q}^{(l)}\in\mathbb{R}^{d}$. Moreover, it is worth noting that the Query component in our architecture does not require the state factorization. Then, the section-adaptive attention without softmax function within the $t$-factor can be expressed as:
$\boldsymbol{a}_{t}^{(l)}=\text{SA}^{t}(\boldsymbol{K}_{t}^{(l)},\boldsymbol{Q}^{(l)})=\frac{{\boldsymbol{K}_{t}^{(l)\top}\boldsymbol{Q}^{(l)}}}{\sqrt{d}},$ (8)
where the achieved $\boldsymbol{a}_{t}^{(l)}\in\mathbb{R}^{n}$ is the section-adaptive attention matrix of the $t$-th factor at the $l$-layer.
Multi-factor Section-adaptive Attention (MFSA). In order to achieve an effective representation fusion of multiple factors, it is essential to effectively integrate the weights allocated to different factors across nodes. Particularly, we recombine a new attention matrix, which is actually formed by combining different section-adaptive attention matrices, and then apply the softmax function on the dimensions of these factors:
$\widetilde{\boldsymbol{A}^{(l)}}=\text{softmax}(\boldsymbol{A}^{(l)}),$ (9)
$\boldsymbol{A}^{(l)}=[\boldsymbol{a}_{1}^{(l)},...,\boldsymbol{a}_{t}^{(l)},...,\boldsymbol{a}_{k}^{(l)}],\;\,t=1,...,k,$ (10)
where $\widetilde{\boldsymbol{A}^{(l)}}\in\mathbb{R}^{n\times k}$ is the resulting matrix by applying the softmax function along the factor’s dimension. Finally, the Multi-factor Section-adaptive Attention (MFSA) mechanism:
[…]
@@ 3.2 Powerformer for Sectional Power System
where $\boldsymbol{V}_{t}^{(l)}$ and $\widetilde{\boldsymbol{A}^{(l)}}$ are obtained from Eq. 5 and 9, $\widetilde{\boldsymbol{A}^{(l)}}_{t}$ denotes $t$-th column associated with $t$-th electrical factor. Symbol $\odot$ denotes the Hadamard product with broadcast mechanism. It is important to note that the left matrix side of the $\odot$ operation is in $\mathbb{R}^{d\times n}$, while the right matrix is in $\mathbb{R}^{n}$. The resulting matrix after the Hadamard product will also reside in the space $\mathbb{R}^{d\times n}$.
3.3 The Implementation of Powerformer

### v4

@@
Abstract.
In this paper, we present a novel transformer architecture tailored for learning robust power system state representations, which strives to optimize power dispatch for the power flow adjustment across different transmission sections. Specifically, our proposed approach, named Powerformer, develops a dedicated section-adaptive attention mechanism, separating itself from the self-attention employed in conventional transformers. This mechanism effectively integrates power system states with transmission section information, which facilitates the development of robust state representations. Furthermore, by considering the graph topology of power system and the electrical attributes of bus nodes, we introduce two customized strategies to further enhance the expressiveness: graph neural network propagation and multi-factor attention mechanism. Extensive evaluations are conducted on three power system scenarios, including the IEEE 118-bus system, a realistic China 300-bus system, and a large-scale European system with 9241 buses, where Powerformer demonstrates its superior performance over several popular baseline methods.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data mining
1. Introduction
Power flow adjustment across the transmission sections refers to manage and regulate the power flow between two system components or entities. As shown in Fig. 1, a transmission section consists of a group of transmission lines with the same color that have the same direction of active power flow and are closely located electrically. In a real-world scenario, the power demand in China’s northern area drastically increases due to the arrival of winter, and the power generation in this area falls insufficient, making it necessary to transfer power from the southern to the northern through the transmission section. To ensure power system security margins, operators commonly employ the total transfer capability of the transmission section to monitor power flows and regulate the operational state (Min and Abur, 2006; Liu and Chu, 2015; Jiang et al., 2022). Specifically, overloading a critical transmission section poses a serious threat to the system, potentially triggering a chain reaction of failures that may cause severe cascading blackouts. Hence, the efficient management of power flow at transmission sections is paramount, serving as a fundamental pillar to ensure the stability, resilience, and reliability of the power system.
Figure 1. Illustration of 10 transmission sections $\{\Phi_{1},...,\Phi_{10}\}$ in the IEEE 118-bus system, where each section is represented by a set of transmission lines that share the same color and are closely located electrically. An illustrative section $\Phi_{7}$ is the set of purple lines, i.e., $\Phi_{7}=\,${bus33-bus37, bus19-bus34, bus30-bus38, bus23-bus24}, that partitions the system into distinct left and right components.
To address power flow adjustment across different transmission sections, power dispatch is an effective strategy that enables precise finetuning of generation outputs to optimize power flow distribution. Traditional dispatch methods, which rely on the system model, can be primarily categorized into two distinct classes: optimization-based (Capitanescu et al., 2011) and sensitivity-based (Dutta and Singh, 2008) techniques. Optimization-based methods utilize numerical techniques to transform the dispatch problem into a constrained programming problem, while sensitivity-based methods apply iterative calculations to find candidate generators by calculating sensitivity indices. However, these model-based methods often encounter the computational burdens and struggle to effectively control unpredictable system coupling within a limited time. Thus, traditional methods are unsuitable for effectively managing real-time control in a complex and large-scale power system.
Recently, deep reinforcement learning (DRL) methods including DQN (Mnih et al., 2015), PPO (Schulman et al., 2017), and A2C (Mnih et al., 2016) have gained attention for their potential in addressing various control challenges within power systems, such as voltage control (Mu et al., 2024; Yang et al., 2020; Cao et al., 2022), economic dispatch (Hu et al., 2024; Dai et al., 2020; Li et al., 2021), and power flow adjustment (Wu et al., 2024; Gao et al., 2021; Wang et al., 2022). For the specific issue of power flow adjustment, these data-driven approaches effectively utilizes the robust capabilities of the pertained deep neural networks, enabling the direct learning of response actions in an end-to-end fashion, effectively tackling the high-dimensional control problems and providing a powerful solution for power flow adjustment. However, among the existing DRL-based methods (Wu et al., 2024; Gao et al., 2021; Wang et al., 2022), there are two evident limitations in the state representations for transmission section power flow adjustment:
  1. (1)
    The absence of efficient awareness regarding transmission sections, which hinders the DRL agent from making optimal action for distinct section adjustment.
  2. (2)
    The high coupled electrical factors and the ignored topology structure of power system, ignificantly affect the extraction of the state characteristics.
In this paper, we strive to develop an advanced transformer architecture for learning robust power system state representations, enabling the DRL agent to perform precise power dispatch strategies that stabilize the power system under various section settings. To tackle the first limitation, we develop a section-adaptive attention mechanism to enable an efficient amalgamation of section information with the power system state, leveraging the comprehensive capabilities of the transformer architecture (Vaswani et al., 2017; Zhu et al., 2023; Zhang et al., 2022; Ding et al., 2024) for the specific sectional system analysis. For the second limitation, we begin by analyzing the graph topology of the power system, and employing graph neural networks (Jing, 2023; Li et al., 2023; Zhang et al., 2023) to propagate features for embedding structural information into state representations. Furthermore, we introduce a multi-factor attention mechanism to efficiently fuse each state factor, ensuring comprehensive and accurate extraction of system states. Finally, we utilize the proposed Powerformer, integrating the Dueling DQN (Wang et al., 2016) framework to achieve precise and efficient control of power flow across transmission sections, thereby leading to improvements in the overall system performance.
To the best of our knowledge, this is the first Transformer architecture that attempts to amalgamate the section information with state representation for transmission section power flow adjustment. Our main contributions can be summarized as follows:
  • Extensive evaluation on three distinct power system cases, including the IEEE 118-bus system, a realistic China 300-bus system, and a large-scale 9241-bus European system, demonstrates our method’s superiority with significantly improved performance.
2. Preliminary
2.1. Problem Formulation
Transmission section power flow adjustment involves the regulation and control of electrical power flow through the specific section or interface of a power transmission network. In a specific power system, there exist $N$ transmission sections $\{\Phi_{1},\Phi_{2},...,\Phi_{N}\}$, where each section $\Phi_{i}$ consists of several transmission lines that not only share the same direction of active power flow but also maintain close electrical proximity to one another (Liu et al., 2023). To accurately determine the total power flow of a transmission section $\Phi_{i}$, the combined power contributions are aggregated and meticulously calculated:
(1) $\mathcal{P}_{\Phi_{i}}=\sum_{j\in\Phi_{i}}\mathcal{P}_{j}\quad\,{s.t.}\;\mathcal{P}_{\text{min},\Phi_{i}}\leq\mathcal{P}_{\Phi_{i}}\leq\mathcal{P}_{\text{max},\Phi_{i}},$
where $\mathcal{P}_{\Phi_{i}}$ and $\mathcal{Q}_{\Phi_{i}}$ represent the total active and reactive power flow in the $i$-th transmission section, and $\mathcal{P}_{j}$ and $\mathcal{Q}_{j}$ denote the active and reactive power flow in the $j$-th transmission line within the section $\Phi_{i}$. $\mathcal{P}_{\text{min},\Phi_{i}}$, $\mathcal{P}_{\text{max},\Phi_{i}}$, $\mathcal{Q}_{\text{min},\Phi_{i}}$, $\mathcal{Q}_{\text{max},\Phi_{i}}$ are the minimum and maximum allowable active and reactive power flow in the $i$-th transmission section. These limits are essential for ensuring the power system operates safely and efficiently, maintaining power flow within each transmission section at levels that prevent overloading and stability issues. To address this problem, power dispatch via DRL is a commonly employed strategy for regulating power system flow and maintaining safety margin constraints.
2.2. Sectional Power System Network
The power system topology, being non-Euclidean (Chen et al., 2020, 2023c), is typically represented as a graph structure (Chen et al., 2023b, d, a). Therefore, we organize power system data in the form of graph-structured data for learning representations. In a power system graph network, nodes represent buses and edges symbolize transmission lines. Specifically, a power system network with section information can be defined as:
Definition 1 (Sectional Power System Network) A sectional power system network is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{H},\bm{Z}_{\Phi})$, where $\mathcal{V}=\{v_{1},...,v_{n}\}$ and =$\mathcal{E}=\{e_{1},...,e_{m}\}$ represent the sets of $n$ nodes and $m$ edges, respectively. $\bm{H}=[\bm{h}_{1},...,\bm{h}_{n}]\in\mathbb{R}^{d\times n}$ denotes the state feature matrix, where $\bm{h}_{i}$ denotes the $i$-th node. $\bm{Z}_{\Phi}$ is the representation of transmission section $\Phi$. Specifically, nodes’ features are initialized with active power, reactive power, voltage magnitude, and phase angle. $\bm{Z}_{\Phi}$ is initialized using the active power on the $i$-th section.
To illustrate a sectional power system network, Fig. 1 provides an example with 10 sections in the IEEE 118-case system. In this case, transmission section $\Phi_{7}$ is represented by the purple lines, which divides the system into left and right components or entities.
Figure 2. The illustration of the Transformer and Powerformer architectures. (a) Transformer architecture with the multi-head self-attention mechanism (MHA). (b) Powerformer architecture with the multi-factor section-adaptive attention mechanism (MFSA).
2.3. Transformer Architecture
Numerous Pre-trained Language Models (PLMs) employ a multi-layer Transformer architecture (Vaswani et al., 2017) for the purpose of encoding text representations. In each layer of the Transformer, a multi-head self-attention mechanism is utilized to generate contextualized representations for individual text tokens. Let $\bm{H}^{(l)}=[\bm{h}_{1}^{(l)},...,\bm{h}_{n}^{(l)}]$ denote the output feature of the $l$-th Transformer layer, where $\bm{h}_{i}^{(l)}\in\mathbb{R}^{d\times 1}$ is the hidden representation of the text token at position $i$. Subsequently, in the $(l+1)$-th Transformer layer, the multi-head self-attention (MHA) is calculated as follows:
@@ 2.3. Transformer Architecture
(3) $\text{head}^{t}(\bm{H}^{(l)})=\bm{V}_{t}^{(l)}\cdot\text{softmax}\left(\frac{{\bm{K}_{t}^{(l)\top}\bm{Q}_{t}^{(l)}}}{\sqrt{d}}\right),$
(4) $\bm{Q}_{t}^{(l)}=\bm{W}_{Q,t}^{(l)}\bm{H}^{(l)},\;\bm{K}_{t}^{(l)}=\bm{W}_{K,t}^{(l)}\bm{H}^{(l)},\;\bm{V}_{t}^{(l)}=\bm{W}_{V,t}^{(l)}\bm{H}^{(l)},$
@@
3. Method
3.1. Overall Architecture
An overview of the Powerformer architecture is presented in Fig. 2b, illustrating the crucial components for obtaining robust representations of sectional power system networks, such as State Factorization, Graph Neural Network (GNN) propagation, and Section-adaptive Attention, etc. Different from the conventional Transformer, our architecture utilizes the set $\{\bm{H},\bm{Z}_{\Phi},\mathcal{E}\}$ as input, which consists of the state feature matrix, section power flow, and graph structure information. The section-specific information $\bm{Z}_{\Phi}$ is utilized to learn Query representation, enabling the development of section-adaptive attention. To address the specific electrical factors on buses, we disentangle the system state representation into $k$ distinct factors to prevent feature coupling during propagation, and then incorporate the section-adaptive attention matrices to establish the multi-factor section-adaptive attention mechanism. Furthermore, to enhance the expressiveness of Key, Value propagation, we utilize the popular GNN architecture to embed the graph structure information $\mathcal{E}$ into state representations.
3.2. Powerformer for Sectional Power System
If the state matrix is directly inputted into the propagation layer, it will result in the coupling of different factors, causing the redundancy in the learned features and ultimately diminishing the expressive power. Thus, we begin by disentangling the input state feature into $k=4$ factors (i.e., active power, reactive power, voltage magnitude, and phase angle), with the $l$-th layer’s representation of the $t$-th factor denoted by $\bm{H}_{t}^{(l)}\in\mathbb{R}^{d\times n}$.
(5) $\bm{K}_{t}^{(l)}=\mathcal{F}_{K,t}^{(l)}(\bm{H}_{t}^{(l)},\mathcal{E}),\;\bm{V}_{t}^{(l)}=\mathcal{F}_{V,t}^{(l)}(\bm{H}_{t}^{(l)},\mathcal{E}),$
where $\mathcal{F}_{K,t}^{(l)},\mathcal{F}_{V,t}^{(l)}$ represent the GNN propagation to update Key and Value: $\bm{K}_{t}^{(l)},\bm{V}_{t}^{(l)}\in\mathbb{R}^{d\times n}$ of the $t$-th factor. Specifically, this propagation is implemented by Graph Isomorphism Network (GIN) (Xu et al., 2019):
(6) $\bm{h}_{v,t}^{(l)}={\rm MLP}^{(l)}((1+\epsilon^{(l)})\cdot\bm{h}_{v,t}^{(l-1)}+\sum\nolimits_{u\in\mathcal{N}(v)}\bm{h}_{u,t}^{(l-1)}),$
where $\bm{h}_{v,t}^{(l)}$ represents the feature vector of node $v$ of the $t$-th factor at the $l$-th layer. $\mathcal{N}(v)$ refers to the neighborhood set of node $v$ based on the edge set $\mathcal{E}$. MLP is the Multi-layer Perceptron network.
Section-adaptive Attention (SA). For section-adaptive representations, we first utilize the power flow information (i.e., active power) of the transmission lines on the current section to encode the section, and set the power flow information that is not within the section to zero. The resulting initialization representation is $\bm{Z}_{\Phi}$. Then, we use a MLP network to learn the Query component:
(7) $\bm{Q}^{(l)}=\mathcal{M}_{Q}^{(l)}(\bm{Z}_{\Phi}^{(l)}),$
where $\mathcal{M}_{Q}^{(l)}$ is the employed MLP to update section representation. The updated Query is $\bm{Q}^{(l)}\in\mathbb{R}^{d\times 1}$. Moreover, it is worth noting that the Query component in our architecture does not perform the state factorization. Then, the section-adaptive attention without softmax function within the $t$-factor can be expressed as:
(8) $\bm{a}_{t}^{(l)}=\text{SA}^{t}(\bm{K}_{t}^{(l)},\bm{Q}^{(l)})=\frac{{\bm{K}_{t}^{(l)\top}\bm{Q}^{(l)}}}{\sqrt{d}},$
where the achieved $\bm{a}_{t}^{(l)}\in\mathbb{R}^{n\times 1}$ is the section-adaptive attention matrix of the $t$-th factor at the $l$-layer.
Multi-factor Section-adaptive Attention (MFSA). In order to achieve an effective representation fusion of multiple factors, it is essential to effectively integrate the weights allocated to different factors across nodes. Particularly, we recombine a new attention matrix, which is actually formed by combining different section-adaptive attention (Eq.8) matrices, and then apply the softmax function on the dimensions of these factors:
(9) $\widetilde{\bm{A}^{(l)}}=\text{softmax}(\bm{A}^{(l)}),$
(10) $\bm{A}^{(l)}=[\bm{a}_{1}^{(l)},...,\bm{a}_{t}^{(l)},...,\bm{a}_{k}^{(l)}],\;\,t=1,...,k,$
where $\widetilde{\bm{A}^{(l)}}\in\mathbb{R}^{n\times k}$ is the resulting matrix by applying the softmax function along the factor’s dimension. Finally, the Multi-factor Section-adaptive Attention (MFSA) mechanism:
@@ 3.2. Powerformer for Sectional Power System
where $\bm{V}_{t}^{(l)}$ and $\widetilde{\bm{A}^{(l)}}$ are obtained from Eq. 5 and 9, $\widetilde{\bm{A}^{(l)}}_{t}\in\mathbb{R}^{n\times 1}$ the $t$-th column associated with $t$-th electrical factor. Symbol $\odot$ denotes the Hadamard product with broadcast mechanism. It is important to note that the left matrix side of the $\odot$ operation is in $\mathbb{R}^{d\times n}$, while the right matrix is in $\mathbb{R}^{n\times 1}$. The resulting matrix after the Hadamard product will also reside in the space $\mathbb{R}^{d\times n}$.
3.3. The Implementation of Powerformer